OpenStudy (anonymous):

34-5(2+p)=p+6(4-p) A) 0 B)no solution C) 1 D) all real numbers PLZ HELP

OpenStudy (uri):

34-10-5p=p+24-6p 34-10-24=-6p+p+5p 34-34=-6p+6p both will be cancelled...so B)No solution.

OpenStudy (anonymous):

thx

OpenStudy (uri):

yw!